# Flatfish Disease Detection Based on Part Segmentation Approach and Disease Image Generation

### v1

@@ 1 Introduction
To achieve this, we focus only on the visual features of diseases. Therefore, we divide the flatfish into three different body parts (i.e., head, body, and fins) and handle them separately. To divide the flatfish, we propose effective three-part segmentation methods in Section 3. Before the part segmentation, all flatfish images are horizontally aligned through the proposed steps. Then, several critical points are estimated from the aligned flatfish images to determine the three parts: head, body, and fins as shown in Figure. 5. The proposed segmentation methods do not require additional annotations of the pixel labels or burdensome deep learning algorithms [8]. After dividing the parts of the flatfish, a flatfish disease detector is trained separately according to the part with the disease. The separate training of disease detectors for each dividing part resulted in a 7% improvement in performance compared to detecting them as a single disease class. Furthermore, by utilizing the proposed image-generation method for data augmentation, there was a 12% improvement in performance compared to its baseline framework.
@@ 3 Part Segmentation Approach for Disease Detection
The symptoms of flatfish diseases can be displayed anywhere on the body. The appearance of disease symptoms is more related to the body part of the fish rather than the types of diseases as depicted in Figure 1. The goal of our study is not to classify the disease types, but to detect the diseases immediately whether any symptoms have been visually exposed on the flatfish. To achieve this, we focus only on the visual features of diseases. Therefore, this section proposes dividing the flatfish into three different parts (i.e., head, body, fins) to evaluate them separately (Figure. 2).
@@ 3.1 Fish Localization and Horizontal Alignment
An olive flounder detector was trained to localize fish in the images by automatically finding the position and size of the fish in the image. The detection result consists of a four-dimensional vector $\mathbf{D}_{n}=\left(x_{n},y_{n},w_{n},h_{n}\right)$, where $\left(x,y\right)$ is the center position, $\left(w,h\right)$ denotes the width and height of the fish, and $n$ represents the index of the detection, respectively. Any object detectors, such as R-CNN [26], or the YOLO series [27, 28, 29] detectors, can be used for fish localization. Rectangular bounding boxes (bbox) $(\mathbf{D}_{n})$ of the detected fish are obtained, but those rough bbox still include many background areas due to tilted fish directions, as illustrated in Figure. 3 (a)–(c). A possible solution is to manually align the fish along the horizontal axis when taking images (e.g., Figure. 3 (d)). However, a wide variety of fish shapes exist, and aligning live and swimming fish is difficult.
@@
where $m_{x}$ and $m_{y}$ are the means of the $x$ and $y$ values, respectively. Based on PCA, two eigenvalues and eigenvectors are calculated from the covariance matrix $\mathbf{\Sigma}$, and the eigenvector $\mathbf{v}_{\lambda}$ corresponding to a larger eigenvalue $\lambda$ is selected. Then, the eigenvector $\mathbf{v}_{\lambda}$ depicted by the cyan line () in Figure. 3 reflects the rotation of the fish.
By rotating the eigenvector $\mathbf{v}_{\lambda}$ to lie on the horizontal line, the fish are aligned regardless of their initial detection results as in Figure. 3(e). The proposed fish alignment method works well because the flatfishes almost like a two-dimensional object on the floor. Compared to the manual alignment in Figure. 3(d), the proposed method performs accurately aligns diverse flatfish horizontally.
@@ 3.2 Fish Part Segmentation for Disease Detection
In the fisher biology field, fish parts are typically categorized into three main sections: head, body, and tail. According to the report [30], the head includes from the snout to the gill, the body expends from the gill to the anus, and the tail covers from the anus to the front of the caudal fin, as illustrated in Figure.4 (a). In this study, we observed that fish parts can be categorized differently according to the types of their diseases. Based on a careful examination of the diseases in our dataset, we newly categorized fish parts as follows.
@@
Owing to the horizontal alignment proposed in Section 3.1, the tail of the flatfish can be found with the following steps. Initially, a set of fish boundary pixels $b\left({C}_{n}\right)$ from the fish foreground points ${C}_{n}$ are determined. These boundary pixels $b\left({C}_{n}\right)$ are then divided into two subsets: the upper and the lower boundaries, as illustrated in Figure. 5. For the upper boundary, the local minimum point $\mathbf{p}^{(tail)}_{up}$ is estimated, and for the lower boundary,the local maximum point $\mathbf{p}^{(tail)}_{low}$ is estimated. Then, the center position of the tail can be defined by $\mathbf{p}^{(tail)}_{c}=\frac{\mathbf{p}^{(tail)}_{up}+\mathbf{p}^{(tail)}_{low}}{2}$. These points determine the position of the tail and the tail length $L^{(tail)}$ can be measured by calculating the distance between $\mathbf{p}^{(tail)}_{c}$ and the end of the tail position as shown in Figure. 5.
Once the tails are found, the head part can be estimated by fitting an ellipse around the gill of the fish. Thus, the center of the head is set to $\mathbf{p}^{(head)}_{c}$ based on the position of the snout and tail length $L^{(tail)}$. Then, the upper and lower positions of the head $\mathbf{p}^{(head)}_{up},\mathbf{p}^{(head)}_{low}$ can be determined along with the vertical direction from $\mathbf{p}^{(head)}_{c}$. Finally, an ellipse can be fit with two axis lengths: the major axis, defined as the line between two head points $\mathbf{p}^{(head)}_{up}$ and $\mathbf{p}^{(head)}_{low}$, and the minor axis, defined as the tail length $L^{(tail)}$. Similarly, the upper and lower fins are defined based on the ellipsoid fitting between the obtained points and tail length. The overall processes are depicted in Figure. 5. Based on the ellipses and points, the three parts of the fish can be separated. Then, three different disease detectors are trained for these parts. This approach enables effectively evaluating different disease appearances.
@@
Table 1: Number of diseases by body part and side.
@@ 4.1 Real Flatfish Disease Images: FlatIMG
Several diseased flatfish images in FlatIMG are shown in Figure. 1. The disease shapes are diverse and can occur anywhere on the fish. In addition, the heatmaps of the diseases are plotted in Figure. 6. They show where the diseases occur on the ocular-side and blind-sides of the flatfish at a glance. Based on the heatmaps, we noted that diseases mainly occur around gills and fins on both sides. These diseases rarely occur on the body of the blind-side (i.e., the stomach) compared to ocular-side.
@@ 4.2 Flatfish Disease Image Generation
The proposed augmentation framework aimed to enrich the dataset by generating natural-looking fish images with diseases that closely mimic real data. The framework consists of two main methods: 1) disease image patch generation; and 2) disease harmonization. The overall framework is described in Figure. 7.
@@ 4.2.1 Disease Patch Generation
For the training data, the diseased patches were cropped with a sufficient margin to include the whole boundaries of the disease pattern. Then, an image generator was trained using distributions of the disease patches. The generator is represented by $p\left(\mathbf{I}|\mathbf{z}\right)$, where the conditional variable $\mathbf{z}$ denotes an input noise vector, and $p\left(\mathbf{I}|\mathbf{z}\right)$ denotes the generated distribution of the disease image given $\mathbf{z}$. Based on the generator, $p\left(\mathbf{I}|\mathbf{z}\right)$, new disease patches can be created by manipulating the input noise vectors $\mathbf{z}$ as shown in Figure. 8. As a result, a set of generated disease image patches is denoted by
@@ 4.2.2 Disease Harmonization
Figure. 9 illustrates the results of image harmonization. The first row shows the mixed images without image harmonization. They simply replaced the area of the target image with the values of the foreground patch. The color of the mixed area is quite unnatural with its surrounding areas. In contrast, the mixed results based on image harmonization are depicted in the second row of Figure. 9, and harmonize well with the color distributions of the target images. However, the boundaries of the source areas are still unnatural due to the large pixel value differences around the boundaries. To handle this problem, we further perform blurring the boundary of foreground patches to make more naturally harmonized images. The third row of Figure. 9 shows the final results of the harmonization including boundary blurring. As in the figure, the unnaturalness of disease boundaries is mitigated resulting in natural-looking diseased flatfish images.
@@ 5.2 Diseased Flatfish Generation Results
Examples of generated diseased flatfish images are shown in Figure. 11, and $4,773$ different images were generated with $462$ generated fish diseases and $59$ non-diseased fishes. The overall quality of the generated images is natural. Note that the proposed framework can perform more diverse diseased fish image generation when it feeds additional source disease and target images. Compared to the flatfish dataset, the dataset was augmented to about $9.3$ times the size using the generated images. We expect that we can fairly validate the proposed disease detection methods in the following section by using the augmented flatfish images.
@@
Figure. 10 shows the distributions of the training patches and the generated patches using the trained model. The generated disease patches are similar to the training patches but not identical. This result supports that we can generate lots of new disease samples from the model and build new diseased flatfish images. For the target images, $59$ disease-free flatfish images were collected from the web. Then, the source and target image pairs were randomly matched, setting random locations and sizes for the target fish images. For example, the location of the diseased patch should not escape the foreground area of the fish. The size of the disease patches is determined within a reasonable range of 80%–120% compared to its original size.
@@ 5.3 Fish Part Segmentation
A method for segmenting fish into three regions (head, fins, and body) was proposed to enhance the detector performance (Section 3.2). The proposed method takes fish images with backgrounds as input, removes the background, and segments the images into three parts. Figure. 12. illustrates the part segmentation results for the fish. Despite shape variations among individual fish, we properly detected feature points of fish heads and tails ($\mathbf{p}^{(head)}_{up},\mathbf{p}^{(head)}_{low},\mathbf{p}^{(tail)}_{up},\mathbf{p}^{(tail)}_{low}$) and can divide the fish into three parts. Based on the feature points, the ellipses were fitted.
The proposed method was also adopted for the flash salmon disease detection task. Salmon have a thinner body than flatfish; thus, the ellipses were determined slightly differently. The ellipses of salmon were generated by connecting a mean-point between $\mathbf{p}^{(tail)}_{up}$ and $\mathbf{p}^{(tail)}_{low}$, and between $\mathbf{p}^{(head)}_{up}$ and $\mathbf{p}^{(head)}_{low}$. Figure. 12 (b) shows several examples of fish-part segmentation of salmon. These experimental results demonstrate that the proposed technique is not only applicable to flatfish but can also be sufficiently applied to other types of fish.
@@ 5.4 Fish Disease Detection Performance
Figure. 13 illustrates the detection results of various methods. The SSD almost failed to detect any diseases. The Faster-RCNN showed better detection performance than the SSD, but it often missed the diseases. The YOLOv5 exhibits higher detection performance, but it showed several false positives in the background, and could not handle various appearances of diseases. The YOLOv8 may seem to have a similar performance to YOLOv5, but looking at the fourth and last detection results, it can be seen that it detected various forms of diseases. Our proposed method (Ours - DA + PartSeg) improved the performance of its baseline (YOLOv8) and robustly detected all diseases that other detectors missed. In particular, it reduced the false positives in the background and detected various types of disease appearances.

### v2

# Flatfish Lesion Detection Based on Part Segmentation Approach and Lesion Image Generation

@@ 1 Introduction
To achieve this, we focus only on the visual features of diseases. Therefore, we divide the flatfish into three different body parts (i.e., head, body, and fins) and handle them separately. To divide the flatfish, we propose effective three-part segmentation methods in Section 3. Before the part segmentation, all flatfish images are horizontally aligned through the proposed steps. Then, several critical points are estimated from the aligned flatfish images to determine the three parts: head, body, and fins as shown in FIGURE 5. The proposed segmentation methods do not require additional annotations of the pixel labels or burdensome deep learning algorithms [8]. After dividing the parts of the flatfish, a flatfish disease detector is trained separately according to the part with the disease. The separate training of disease detectors for each dividing part resulted in a 7% improvement in performance compared to detecting them as a single disease class. Furthermore, by utilizing the proposed image-generation method for data augmentation, there was a 12% improvement in performance compared to its baseline framework.
@@ 3 Part Segmentation Approach for Disease Detection
The symptoms of flatfish diseases can be displayed anywhere on the body. The appearance of disease symptoms is more related to the body part of the fish rather than the types of diseases as depicted in FIGURE 1. The goal of our study is not to classify the disease types, but to detect the diseases immediately whether any symptoms have been visually exposed on the flatfish. To achieve this, we focus only on the visual features of diseases. Therefore, this section proposes dividing the flatfish into three different parts (i.e., head, body, fins) to evaluate them separately (FIGURE 2).
@@ 3.1 Fish Localization and Horizontal Alignment
An olive flounder detector was trained to localize fish in the images by automatically finding the position and size of the fish in the image. The detection result consists of a four-dimensional vector $\mathbf{D}_{n}=\left(x_{n},y_{n},w_{n},h_{n}\right)$, where $\left(x,y\right)$ is the center position, $\left(w,h\right)$ denotes the width and height of the fish, and $n$ represents the index of the detection, respectively. Any object detectors, such as R-CNN [26], or the YOLO series [27, 28, 29] detectors, can be used for fish localization. Rectangular bounding boxes (bbox) $(\mathbf{D}_{n})$ of the detected fish are obtained, but those rough bbox still include many background areas due to tilted fish directions, as illustrated in FIGURE 3 (a)–(c). A possible solution is to manually align the fish along the horizontal axis when taking images (e.g., FIGURE 3 (d)). However, a wide variety of fish shapes exist, and aligning live and swimming fish is difficult.
@@
where $m_{x}$ and $m_{y}$ are the means of the $x$ and $y$ values, respectively. Based on PCA, two eigenvalues and eigenvectors are calculated from the covariance matrix $\mathbf{\Sigma}$, and the eigenvector $\mathbf{v}_{\lambda}$ corresponding to a larger eigenvalue $\lambda$ is selected. Then, the eigenvector $\mathbf{v}_{\lambda}$ depicted by the cyan line () in FIGURE 3 reflects the rotation of the fish.
By rotating the eigenvector $\mathbf{v}_{\lambda}$ to lie on the horizontal line, the fish are aligned regardless of their initial detection results as in FIGURE 3(e). The proposed fish alignment method works well because the flatfishes almost like a two-dimensional object on the floor. Compared to the manual alignment in FIGURE 3(d), the proposed method performs accurately aligns diverse flatfish horizontally.
@@ 3.2 Fish Part Segmentation for Disease Detection
In the fisher biology field, fish parts are typically categorized into three main sections: head, body, and tail. According to the report [30], the head includes from the snout to the gill, the body expends from the gill to the anus, and the tail covers from the anus to the front of the caudal fin, as illustrated in FIGURE4 (a). In this study, we observed that fish parts can be categorized differently according to the types of their diseases. Based on a careful examination of the diseases in our dataset, we newly categorized fish parts as follows.
@@
Owing to the horizontal alignment proposed in Section 3.1, the tail of the flatfish can be found with the following steps. Initially, a set of fish boundary pixels $b\left({C}_{n}\right)$ from the fish foreground points ${C}_{n}$ are determined. These boundary pixels $b\left({C}_{n}\right)$ are then divided into two subsets: the upper and the lower boundaries, as illustrated in FIGURE 5. For the upper boundary, the local minimum point $\mathbf{p}^{(tail)}_{up}$ is estimated, and for the lower boundary,the local maximum point $\mathbf{p}^{(tail)}_{low}$ is estimated. Then, the center position of the tail can be defined by $\mathbf{p}^{(tail)}_{c}=\frac{\mathbf{p}^{(tail)}_{up}+\mathbf{p}^{(tail)}_{low}}{2}$. These points determine the position of the tail and the tail length $L^{(tail)}$ can be measured by calculating the distance between $\mathbf{p}^{(tail)}_{c}$ and the end of the tail position as shown in FIGURE 5.
Once the tails are found, the head part can be estimated by fitting an ellipse around the gill of the fish. Thus, the center of the head is set to $\mathbf{p}^{(head)}_{c}$ based on the position of the snout and tail length $L^{(tail)}$. Then, the upper and lower positions of the head $\mathbf{p}^{(head)}_{up},\mathbf{p}^{(head)}_{low}$ can be determined along with the vertical direction from $\mathbf{p}^{(head)}_{c}$. Finally, an ellipse can be fit with two axis lengths: the major axis, defined as the line between two head points $\mathbf{p}^{(head)}_{up}$ and $\mathbf{p}^{(head)}_{low}$, and the minor axis, defined as the tail length $L^{(tail)}$. Similarly, the upper and lower fins are defined based on the ellipsoid fitting between the obtained points and tail length. The overall processes are depicted in FIGURE 5. Based on the ellipses and points, the three parts of the fish can be separated. Then, three different disease detectors are trained for these parts. This approach enables effectively evaluating different disease appearances.
@@
Table 1: Number of diseases by body part and side.
@@
Table 2: Number of diseases by pathogen type.
@@ 4.1 Real Flatfish Disease Images: FlatIMG
Several diseased flatfish images in FlatIMG are shown in FIGURE 1. The disease shapes are diverse and can occur anywhere on the fish. In addition, the heatmaps of the diseases are plotted in FIGURE 6. They show where the diseases occur on the ocular-side and blind-sides of the flatfish at a glance. Based on the heatmaps, we noted that diseases mainly occur around gills and fins on both sides. These diseases rarely occur on the body of the blind-side (i.e., the stomach) compared to ocular-side.
@@ 4.2 Flatfish Disease Image Generation
The proposed augmentation framework aimed to enrich the dataset by generating natural-looking fish images with diseases that closely mimic real data. The framework consists of two main methods: 1) disease image patch generation; and 2) disease harmonization. The overall framework is described in FIGURE 7.
@@ 4.2.1 Disease Patch Generation
For the training data, the diseased patches were cropped with a sufficient margin to include the whole boundaries of the disease pattern. Then, an image generator was trained using distributions of the disease patches. The generator is represented by $p\left(\mathbf{I}|\mathbf{z}\right)$, where the conditional variable $\mathbf{z}$ denotes an input noise vector, and $p\left(\mathbf{I}|\mathbf{z}\right)$ denotes the generated distribution of the disease image given $\mathbf{z}$. Based on the generator, $p\left(\mathbf{I}|\mathbf{z}\right)$, new disease patches can be created by manipulating the input noise vectors $\mathbf{z}$ as shown in FIGURE 8. As a result, a set of generated disease image patches is denoted by
@@ 4.2.2 Disease Harmonization
FIGURE 9 illustrates the results of image harmonization. The first row shows the mixed images without image harmonization. They simply replaced the area of the target image with the values of the foreground patch. The color of the mixed area is quite unnatural with its surrounding areas. In contrast, the mixed results based on image harmonization are depicted in the second row of FIGURE 9, and harmonize well with the color distributions of the target images. However, the boundaries of the source areas are still unnatural due to the large pixel value differences around the boundaries. To handle this problem, we further perform blurring the boundary of foreground patches to make more naturally harmonized images. The third row of FIGURE 9 shows the final results of the harmonization including boundary blurring. As in the figure, the unnaturalness of disease boundaries is mitigated resulting in natural-looking diseased flatfish images.
@@ 5.2 Diseased Flatfish Generation Results
Examples of generated diseased flatfish images are shown in FIGURE 11, and $4,773$ different images were generated with $462$ generated fish diseases and $59$ non-diseased fishes. The overall quality of the generated images is natural. Note that the proposed framework can perform more diverse diseased fish image generation when it feeds additional source disease and target images. Compared to the flatfish dataset, the dataset was augmented to about $9.3$ times the size using the generated images. We expect that we can fairly validate the proposed disease detection methods in the following section by using the augmented flatfish images.
@@
FIGURE 10 shows the distributions of the training patches and the generated patches using the trained model. The generated disease patches are similar to the training patches but not identical. This result supports that we can generate lots of new disease samples from the model and build new diseased flatfish images. For the target images, $59$ disease-free flatfish images were collected from the web. Then, the source and target image pairs were randomly matched, setting random locations and sizes for the target fish images. For example, the location of the diseased patch should not escape the foreground area of the fish. The size of the disease patches is determined within a reasonable range of 80%–120% compared to its original size.
@@ 5.3 Fish Part Segmentation
A method for segmenting fish into three regions (head, fins, and body) was proposed to enhance the detector performance (Section 3.2). The proposed method takes fish images with backgrounds as input, removes the background, and segments the images into three parts. FIGURE 12. illustrates the part segmentation results for the fish. Despite shape variations among individual fish, we properly detected feature points of fish heads and tails ($\mathbf{p}^{(head)}_{up},\mathbf{p}^{(head)}_{low},\mathbf{p}^{(tail)}_{up},\mathbf{p}^{(tail)}_{low}$) and can divide the fish into three parts. Based on the feature points, the ellipses were fitted.
The proposed method was also adopted for the flash salmon disease detection task. Salmon have a thinner body than flatfish; thus, the ellipses were determined slightly differently. The ellipses of salmon were generated by connecting a mean-point between $\mathbf{p}^{(tail)}_{up}$ and $\mathbf{p}^{(tail)}_{low}$, and between $\mathbf{p}^{(head)}_{up}$ and $\mathbf{p}^{(head)}_{low}$. FIGURE 12 (b) shows several examples of fish-part segmentation of salmon. These experimental results demonstrate that the proposed technique is not only applicable to flatfish but can also be sufficiently applied to other types of fish.
@@ 5.4 Fish Disease Detection Performance
FIGURE 13 illustrates the detection results of various methods. The SSD almost failed to detect any diseases. The Faster-RCNN showed better detection performance than the SSD, but it often missed the diseases. The YOLOv5 exhibits higher detection performance, but it showed several false positives in the background, and could not handle various appearances of diseases. The YOLOv8 may seem to have a similar performance to YOLOv5, but looking at the fourth and last detection results, it can be seen that it detected various forms of diseases. Our proposed method (Ours - DA + PartSeg) improved the performance of its baseline (YOLOv8) and robustly detected all diseases that other detectors missed. In particular, it reduced the false positives in the background and detected various types of disease appearances.